You measure 33 textbooks weights, and find they have a mean weight of 43 ounces. Assume the population standard deviation is 4.6 ounces. Based on this, construct a 90% confidence interval for the true population mean textbook weight. Give your answers as decimals, to two places License Question 4 of 7 Points possible: 1 This is attempt 1 of 1 .
Answer #1

90% Confidence Interval of true population mean:

[\overline{X} - \tau_{0.95} \frac{\sigma}{\sqrt{n}}, \overline{X} + \tau_{0.95} \frac{\sigma}{\sqrt{n}} ], where, \overline{X} = 43, \sigma = 4.6, \tau_{0.95} = 1.645,

n = 33.

= [43 - 1.3172, 43 + 1.3172] = [41.6828, 44.3172].

Hence, 41.68 < \mu < 44.32. (Ans).